I need to write the equation of the line that passes trhough P(-3,6) and perpendicular to the graph of y=-2/3x-7.
Answer by stanbon(75887)

I need to write the equation of the line that passes trhough P(-3,6) and perpendicular to the graph of y=-2/3x-7.
---
The given line has slope = -2/3
---
Any line perpendicular to it must have slope = 3/2
----------------
Form: y = mx + b
6 = (3/2)(-3)+ b
b = 6+ 9/2
---
b = 10.5
------
Equation:
y = (3/2)x + (21/2)
=======================
Cheers,
Stan H.